OpenStudy (anonymous):

The size of a bicycle is determined by the diameter of the wheel. You want a 22-in bicycle for your birthday. What is the area of one of the wheels?(

OpenStudy (mathmagician):

A=pi*(d/2)^=3.14159*11^2=380.13239

OpenStudy (anonymous):

THANKS